Ring current 6.67e7 particles/s
pressure at S1: 2.3-2.8e-8
Merger, Tape server (WS3): 1574 kB/s, writing ok
Leakage current (WS4): 3.20 uA
DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed
Online monitoring (WS 5): done |